# Some phenomenological aspects of a quantum-corrected Reissner–Nordström black hole: quasi-periodic oscillations, scalar perturbations and thermal fluctuations

### v1

@@ Abstract
In this work, we investigate several phenomenological aspects of a covariant quantum-corrected Reissner–Nordström black hole characterized by the mass $M$, electric charge $Q$, and the quantum correction parameter $\zeta$. We first study the motion of neutral test particles and derive the fundamental orbital and epicyclic frequencies, which are then employed to analyze different quasi–periodic oscillation (QPO) models. Using observational QPO data from stellar–mass, intermediate–mass, and supermassive black hole candidates, we perform a Bayesian parameter estimation through a Markov Chain Monte Carlo (MCMC) analysis and obtain constraints on the black hole parameters. The results show that the presence of the quantum correction significantly affects the location of the QPO radii and the separation between the QPO orbit and the ISCO. We then examine the scalar perturbations by deriving the Schrödinger–like radial equation and the corresponding effective potential. The influence of the parameters $Q$ and $\zeta$ on the perturbation potential and stability of the spacetime is discussed. Furthermore, we compute the greybody factor and the energy emission rate in the high–frequency (geometric–optics) regime, showing how the quantum correction modifies the absorption probability and radiation spectrum. Finally, we study the effect of thermal fluctuations on the black hole entropy and obtain the logarithmic corrections to the Bekenstein–Hawking area law. We show that these corrections become important for small black holes, while for large horizon radius the standard thermodynamic behavior is recovered. Our analysis demonstrates that the quantum correction parameter leaves observable imprints on both dynamical and thermodynamical properties of the spacetime and can be constrained through QPO observations.
@@ I Introduction
Black holes (BHs), predicted by General Relativity (GR), have now been confirmed through various observational means [1, 26]. Their intense gravitational fields make them ideal laboratories for testing diverse gravitational theories. Simultaneously, BHs possess enormous energy. Phenomena such as high-energy radiation and relativistic jets produced by the accretion disks surrounding BHs [27] constitute some of the most active and energetically efficient astrophysical processes in the Universe. It was generally believed that this immense energy was inaccessible to external extraction. This view changed when Penrose, based on the existence of negative-energy orbits within the ergoregion, proposed a mechanism for extracting energy from a rotating BH, later known as the Penrose process [56]. This proposal fundamentally transformed our understanding of BHs, redefining them from passive gravitational endpoints into active media capable of exchanging energy with their surroundings.
@@
This study investigates the dynamics of particles and the resulting QPOs in the spacetime of a quantum-corrected Reissner-Nordström (RN) BH. The paper is organized as follows. In Sec. II, we first provide a brief review on the quantum-corrected RN BH and analyze its horizon structure. We then investigate the motion of test particles around this BH in Sec. III. Fundamental frequencies were discussed in Sec. IV. The scalar perturbations of the BH and greybody factors in Sec. V. The energy emission and thermal fluctuations are discussed in Secs. VI and VII. Finally, the conclusions in Sec. VIII.
@@ III Dynamics of Neutral particles
In this section, we investigate the dynamics of test particles around the quantum-corrected RN BH. Using the Hamiltonian formalism, we derive the effective potential, from which the specific energy and angular momentum of particles in circular orbits can be obtained by imposing the circular orbit conditions. Furthermore, we analyze the marginally stable circular orbits. The dynamics of neutral particles around BHs in various configurations have been extensively studied in the literature (see [51, 7, 9, 6, 5, 8] and related references therein).
@@ IV Fundamental Frequencies
To study the QPO properties of the QCRN BH, we calculate the fundamental frequencies of massive neutral test particles orbiting the BH. In this work, we adopt one of the simplest and most widely used approaches, where the accretion dynamics is modeled by test particles moving along timelike geodesics in the given spacetime geometry. Despite its simplicity, this model captures the essential physical mechanisms responsible for QPO observables around compact astrophysical objects.
We consider small perturbations around stable circular equatorial orbits. When a particle is slightly displaced from its equilibrium position, its motion can be described by radial and vertical epicyclic oscillations, which to leading order behave as linear harmonic oscillators. This perturbative framework allows us to define the Keplerian, radial, and vertical epicyclic frequencies measured by a distant observer, and to analyze the stability properties of circular orbits in the QCRN spacetime.
@@ IV.2 Harmonic oscillations
Also, as it can be observed, $\Omega_{\theta}=\Omega_{\phi}=\Omega_{K}$. Considering the lapse function of the QCRN spacetime in Eq. (2), one obtains
@@ IV.3 QPO models
In Fig. 6, we illustrate the correlated behavior of the upper and lower frequencies of twin-peaked QPOs in the vicinity of QCRN BHs, together with their Schwarzschild and RN limits. The frequencies are shown for radii ranging from the ISCO out to infinity, and for several representative values of the BH parameters $Q$ and $\zeta$. The diagrams display a set of inclined lines corresponding to specific frequency ratios, namely $\nu_{U}\!:\!\nu_{L}=3\!:\!2$, 4:34\!:\!3, 5:45\!:\!4, and 1:11\!:\!1. These lines indicate the allowed combinations of upper and lower frequencies compatible with each ratio [44, 42, 72, 71]. In particular, the 1:11\!:\!1 line—often referred to as the “graveyard” of twin-peaked QPOs—marks the regime where the two frequencies coincide, so that the system effectively exhibits a single QPO peak.
Figure 6: Relations between the upper and lower frequencies of twin-peaked QPOs in the RP, ER2–ER4, and WD models around QCRN BHs, shown for several values of $Q$ and $\zeta$. The straight lines indicate, from top to bottom, the frequency ratios $\nu_{U}\!:\!\nu_{L}=3\!:\!2$, 4:34\!:\!3, 5:45\!:\!4, and 1:11\!:\!1.
@@ IV.4 BH parameter estimation from QPO data
In this subsection, we aim to constrain the parameters associated with particle motion around the QCRN BH by confronting the model with observational data from three classes of BH candidates: stellar-mass, intermediate-mass, and supermassive systems.
@@
We then employ the Python package emcee [29, 89, 2] to carry out an MCMC analysis and to constrain the parameters describing the motion of test particles around the QCRN BH.
@@
where $\theta_{i}=\{r/M,M/M_{\odot},Q/M,\zeta/M\}$ for the spacetime of the QCRN BH, and $\sigma_{i}$ are their respective variances. Furthermore, to perform our MCMC analysis, we use two distinctive data parts, which result in the likelihood functions
@@
Table 2: The Gaussian priors (including the mean values $\mu$ and the variances $\sigma$) given for the QCRN BH within the considered QPO data sources.
@@
Table 3: Best–fit values and constraints on the QCRN BH parameters inferred from the QPO data listed in Table 1.
@@
Figure 10: Constraints on the QCRN BH parameters obtained from a four–dimensional MCMC analysis using the QPO data sets.
To constrain the parameter set $\{r/M,\,M/M_{\odot},\,Q/M,\,\zeta/M\}$ of the QCRN BH, we perform a MCMC analysis employing the Gaussian priors introduced above, which are based on the observational inputs for the QPO sources listed in Table 1. For each parameter, we draw approximately $10^{5}$ samples from the corresponding prior Gaussian distribution. This procedure enables an efficient exploration of the physically admissible region of the parameter space within the adopted bounds and allows us to determine the best–fit values of the model parameters, which are reported in Table 3.
@@
According to the diagrams, one can infer that in all cases, the MCMC chains converge to well-defined regions in the parameter space $\{r/M,\,M/M_{\odot},\,Q/M,\,\zeta/M\}$, indicating that the QCRN model is capable of providing meaningful constraints from the observational data. The stellar-mass systems XTE J1550-564 and GRO J1655-40 exhibit relatively tight posteriors, especially for $M/M_{\odot}$ and $Q/M$, reflecting the higher precision of their QPO measurements. In contrast, the supermassive source Sgr A* shows broader credible regions, particularly in $\zeta/M$, signaling a larger degree of degeneracy among the model parameters. The intermediate-mass candidate M82-X1 lies between these two regimes, with uncertainties larger than those of the stellar-mass sources but smaller than those of Sgr A*. Despite these differences, all four sources display consistent, well-behaved posterior structures, supporting the robustness of the inferred constraints within the QCRN framework.
V Scalar Perturbations; The Massless Klein-Gordon Equation
In this section, we investigate the dynamics of a massless scalar field around a quantum-corrected charged BH, highlighting how quantum corrections and other BH parameters modify the perturbation potential. A large body of work has been devoted to studying scalar perturbations in BHs, both within the framework of GR and in various modified gravity theories.
The massless scalar field dynamics are governed by the Klein-Gordon equation [18, 41, 70, 34, 46, 14]
@@ V Scalar Perturbations; The Massless Klein-Gordon Equation
$\Bigg[\left(\frac{2M}{r}-\frac{2Q^{2}}{r^{2}}\right)\left\{1+\frac{2\zeta^{2}}{r^{2}}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)\right\}$
@@
(i) $\ell=0$                   (ii) $\ell=1$
Figure 11: Behavior of the scalar perturbation potential of quantum-corrected RN BH for various values of the quantum correction parameter $\zeta$. The rest of the arguments are set as: $Q/M=1$.
(i) $\ell=0$                   (ii) $\ell=1$
Figure 12: 3D contour plot of the perturbations potential as a function $(r/M,\,\zeta/M)$. Here $Q/M=1$
@@
    $V_{\text{s}}(r)$ $=$ $\frac{1}{r^{2}}\left(\frac{2M}{r}-\frac{2Q^{2}}{r^{2}}\right)\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (59)
@@
    $\Bigg[\frac{2M}{r}\Bigl(1+\frac{2\zeta^{2}}{r^{2}}\left(1-\frac{2M}{r}\right)\Bigr)-\frac{2\zeta^{2}}{r^{2}}\left(1-\frac{2M}{r}\right)^{2}\Bigg].$ (60)
@@
V.1 Greybody Factor
We will use the general method of semi-analytical bounds for the greybody factors [49, 86, 10] to compute the greybody factor of the quantum-corrected RN BH spacetime as follows,
@@ V.1 Greybody Factor
$\wp=\frac{1}{2h}\sqrt{\left(\frac{dh\left(r_{\ast}\right)}{dr_{\ast}}\right)^{2}+(\omega^{2}-V_{eff}-h^{2}\left(r_{\ast}\right))^{2}},$ (62)
@@
$T\left(\omega\right)\geq\sec h^{2}\left(\int_{r_{h}}^{+\infty}\frac{V_{eff}}{2\omega}dr_{\ast}\right).$ (63)
@@
VI Energy emission
@@
In the geometric-optics (high-frequency) regime, the absorption cross section oscillates around a constant limiting value $\sigma_{\rm lim}$. Motivated by the fact that the capture of high-energy quanta is governed by null geodesics, one may approximate this limiting cross section by a horizon-scale area, $\sigma_{\rm lim}\approx\pi r_{h}^{2}$ [25], where $r_{h}$ is the event-horizon radius. Within this approximation, the spectral energy emission rate takes the standard blackbody-like form [52]
$\frac{d^{2}\varepsilon}{d\omega\,dt}=\frac{2\pi^{2}\sigma_{\rm lim}}{e^{\omega/T}-1}\,\omega^{3},$ (65)
@@
As the selected space-time is asymptotically flat, the Hawking temperature is given by
$T=\frac{f^{\prime}(r_{h})}{4\pi}=\frac{1}{2\pi r_{h}}\left(\frac{M}{r_{h}}-\frac{Q^{2}}{r_{h}^{2}}\right).$ (66)
@@
$\frac{d^{2}\varepsilon}{d\omega\,dt}=2\pi^{3}\,r_{h}^{2}\,\omega^{3}\left[\exp\left\{2\pi r_{h}\,\omega\,\left(\frac{M}{r_{h}}-\frac{Q^{2}}{r_{h}^{2}}\right)^{-1}\right\}-1\right]^{-1},$ (67)

### v2

@@ Abstract
In this work, we investigate several phenomenological aspects of a covariant quantum-corrected Reissner-Nordström black hole characterized by the mass $M$, electric charge $Q$, and the quantum correction parameter $\zeta$. We first study the motion of neutral test particles and derive the fundamental orbital and epicyclic frequencies, which are then employed to analyze different quasi–periodic oscillation (QPO) models. Using observational QPO data from stellar–mass, intermediate–mass, and supermassive black hole candidates, we perform a Bayesian parameter estimation through a Markov Chain Monte Carlo (MCMC) analysis and obtain constraints on the black hole parameters. The results show that the presence of the quantum correction significantly affects the location of the QPO radii and the separation between the QPO orbit and the ISCO. We then examine the scalar perturbations by deriving the Schrödinger-like radial equation and the corresponding effective potential. The influence of the parameters $Q$ and $\zeta$ on the perturbation potential and stability of the spacetime is discussed. Furthermore, we compute the greybody factor and the energy emission rate in the high-frequency (geometric–optics) regime, showing how the quantum correction modifies the absorption probability and radiation spectrum. Finally, we study the effect of thermal fluctuations on the black hole entropy and obtain the logarithmic corrections to the Bekenstein-Hawking area law. We show that these corrections become important for small black holes, while for large horizon radius the standard thermodynamic behavior is recovered. Our analysis demonstrates that the quantum correction parameter leaves observable imprints on both dynamical and thermodynamical properties of the spacetime and can be constrained through QPO observations.
arXiv:2602.15551 [gr-qc]
@@ I Introduction
Black holes (BHs), predicted by General Relativity (GR), have now been confirmed through various observational means [2, 29, 30]. Their intense gravitational fields make them ideal laboratories for testing diverse gravitational theories. Simultaneously, BHs possess enormous energy. Phenomena such as high-energy radiation and relativistic jets produced by the accretion disks surrounding BHs [31] constitute some of the most active and energetically efficient astrophysical processes in the Universe. It was generally believed that this immense energy was inaccessible to external extraction. This view changed when Penrose, based on the existence of negative-energy orbits within the ergoregion, proposed a mechanism for extracting energy from a rotating BH, later known as the Penrose process [62]. This proposal fundamentally transformed our understanding of BHs, redefining them from passive gravitational endpoints into active media capable of exchanging energy with their surroundings.
@@
This study investigates the dynamics of particles and the resulting QPOs in the spacetime of a quantum-corrected Reissner-Nordström (RN) BH. The paper is organized as follows. In Sec. II, we first provide a brief review on the quantum-corrected RN BH and analyze its horizon structure. We then investigate the motion of test particles around this BH in Sec. III. Fundamental frequencies were discussed in Sec. IV. The scalar perturbations of the BH and greybody factors in Sec. V. The energy emission rate in Sec. VI and thermal fluctuations in Sec.VII. Finally, the conclusions in Sec. VIII.
@@ III Dynamics of Neutral particles
In this section, we investigate the dynamics of test particles around the quantum-corrected RN BH. Using the Hamiltonian formalism, we derive the effective potential, from which the specific energy and angular momentum of particles in circular orbits can be obtained by imposing the circular orbit conditions. Furthermore, we analyze the marginally stable circular orbits. The dynamics of neutral particles around BHs in various configurations have been extensively studied in the literature (see [22, 92, 57, 16, 14, 12, 10] and related references therein).
@@ IV Fundamental Frequencies
To study the quasi-periodic oscillations (QPOs) properties of the quantum-corrected RN BH, we calculate the fundamental frequencies of massive neutral test particles orbiting the BH. In this work, we adopt one of the simplest and most widely used approaches, where the accretion dynamics is modeled by test particles moving along timelike geodesics in the given spacetime geometry. Despite its simplicity, this model captures the essential physical mechanisms responsible for QPOs observables around compact astrophysical objects.
We consider small perturbations around stable circular equatorial orbits. When a particle is slightly displaced from its equilibrium position, its motion can be described by radial and vertical epicyclic oscillations, which to leading order behave as linear harmonic oscillators. This perturbative framework allows us to define the Keplerian, radial, and vertical epicyclic frequencies measured by a distant observer, and to analyze the stability properties of circular orbits in the quantum-corrected RN spacetime. It is worth noting that QPOs for quantum-corrected uncharged BHs have recently been studied in Ref. [98]. Subsequent analyses using Event Horizon Telescope (EHT) and X-ray binary observations, as well as astrophysical constraints from the orbit of the S2 star, were performed in Refs. [15, 94]. In the present work, we extend these investigations to the charged sector by considering a quantum-corrected RN black hole and analyzing its associated QPO properties.
@@ IV.2 Harmonic oscillations
Also, as it can be observed, $\Omega_{\theta}=\Omega_{\phi}=\Omega_{K}$. Considering the lapse function of the quantum-corrected RN spacetime in Eq. (2), one obtains
@@ IV.3 QPO models
In Fig. 6, we illustrate the correlated behavior of the upper and lower frequencies of twin-peaked QPOs in the vicinity of quantum-corrected RN BHs, together with their Schwarzschild and RN limits. The frequencies are shown for radii ranging from the ISCO out to infinity, and for several representative values of the BH parameters $Q$ and $\zeta$. The diagrams display a set of inclined lines corresponding to specific frequency ratios, namely $\nu_{U}\!:\!\nu_{L}=3\!:\!2$, 4:34\!:\!3, 5:45\!:\!4, and 1:11\!:\!1. These lines indicate the allowed combinations of upper and lower frequencies compatible with each ratio [48, 46, 77, 76]. In particular, the 1:11\!:\!1 line—often referred to as the “graveyard” of twin-peaked QPOs—marks the regime where the two frequencies coincide, so that the system effectively exhibits a single QPO peak.
Figure 6: Relations between the upper and lower frequencies of twin-peaked QPOs in the RP, ER2–ER4, and WD models around quantum-corrected RN BHs, shown for several values of $Q$ and $\zeta$. The straight lines indicate, from top to bottom, the frequency ratios $\nu_{U}\!:\!\nu_{L}=3\!:\!2$, 4:34\!:\!3, 5:45\!:\!4, and 1:11\!:\!1.
@@ IV.4 BH parameter estimation from QPO data
In this subsection, we aim to constrain the parameters associated with particle motion around the quantum-corrected RN BH by confronting the model with observational data from three classes of BH candidates: stellar-mass, intermediate-mass, and supermassive systems.
@@
We then employ the Python package emcee [33, 97, 6] to carry out an MCMC analysis and to constrain the parameters describing the motion of test particles around the quantum-corrected RN BH.
@@
where $\theta_{i}=\{r/M,M/M_{\odot},Q/M,\zeta/M\}$ for the spacetime of the quantum-corrected RN BH, and $\sigma_{i}$ are their respective variances. Furthermore, to perform our MCMC analysis, we use two distinctive data parts, which result in the likelihood functions
@@
Table 2: The Gaussian priors (including the mean values $\mu$ and the variances $\sigma$) given for the quantum-corrected RN BH within the considered QPO data sources.
@@
Table 3: Best–fit values and constraints on the quantum-corrected RN BH parameters inferred from the QPO data listed in Table 1.
@@
Figure 10: Constraints on the quantum-corrected RN BH parameters obtained from a four–dimensional MCMC analysis using the QPO data sets.
To constrain the parameter set $\{r/M,\,M/M_{\odot},\,Q/M,\,\zeta/M\}$ of the quantum-corrected RN BH, we perform a MCMC analysis employing the Gaussian priors introduced above, which are based on the observational inputs for the QPO sources listed in Table 1. For each parameter, we draw approximately $10^{5}$ samples from the corresponding prior Gaussian distribution. This procedure enables an efficient exploration of the physically admissible region of the parameter space within the adopted bounds and allows us to determine the best–fit values of the model parameters, which are reported in Table 3.
@@
According to the diagrams, one can infer that in all cases, the MCMC chains converge to well-defined regions in the parameter space $\{r/M,\,M/M_{\odot},\,Q/M,\,\zeta/M\}$, indicating that the quantum-corrected RN model is capable of providing meaningful constraints from the observational data. The stellar-mass systems XTE J1550-564 and GRO J1655-40 exhibit relatively tight posteriors, especially for $M/M_{\odot}$ and $Q/M$, reflecting the higher precision of their QPO measurements. In contrast, the supermassive source Sgr A* shows broader credible regions, particularly in $\zeta/M$, signaling a larger degree of degeneracy among the model parameters. The intermediate-mass candidate M82-X1 lies between these two regimes, with uncertainties larger than those of the stellar-mass sources but smaller than those of Sgr A*. Despite these differences, all four sources display consistent, well-behaved posterior structures, supporting the robustness of the inferred constraints within the quantum-corrected RN framework.
V Scalar Field Perturbations
In this section, we investigate the dynamics of a massless scalar field around a quantum-corrected charged black hole, highlighting how quantum corrections and other black hole parameters modify the perturbation potential. Using this potential, we analyze the greybody factors, which encode the modification of Hawking radiation due to the black hole’s potential barrier, determining the fraction of particles that can escape to infinity. Greybody factors are therefore crucial for understanding the black hole’s energy emission spectrum and its observational signatures. A large body of work has been devoted to studying scalar perturbations of black holes, both within the framework of general relativity and in various modified gravity theories (see [27, 45, 75, 38, 50, 23, 11, 17, 9, 13]).
V.1 The Massless Klein-Gordon Equation
The massless scalar field dynamics are governed by the Klein-Gordon equation [27]
@@ V.1 The Massless Klein-Gordon Equation
$\Bigg[\ell(\ell+1)+\left(\frac{2M}{r}-\frac{2Q^{2}}{r^{2}}\right)\left\{1+\frac{2\zeta^{2}}{r^{2}}\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)\right\}$
@@
(i) $\ell=0$
(ii) $\ell=1$
Figure 11: Behaviors of the scalar perturbation potential of quantum-corrected RN BH for various values of the quantum correction parameter $\zeta$. The rest of the arguments are set as: $Q/M=1$.
@@
    $V_{\text{s}}(r)=\frac{1}{r^{2}}\left(\ell(\ell+1)+\frac{2M}{r}-\frac{2Q^{2}}{r^{2}}\right)\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)$ (59)
@@
    $\Bigg[\ell(\ell+1)+\frac{2M}{r}\left\{1+\frac{2\zeta^{2}}{r^{2}}\left(1-\frac{2M}{r}\right)\right\}-\frac{2\zeta^{2}}{r^{2}}\left(1-\frac{2M}{r}\right)^{2}\Bigg].$ (60)
@@
V.2 Greybody Factor
We will use the general method of semi-analytical bounds for the greybody factors [54, 91, 18, 9, 13] to compute the greybody factor of the quantum-corrected RN BH spacetime as follows,
@@ V.2 Greybody Factor
$\wp=\frac{1}{2h}\sqrt{\left(\frac{dh\left(r_{\ast}\right)}{dr_{\ast}}\right)^{2}+(\omega^{2}-V_{s}-h^{2}\left(r_{\ast}\right))^{2}},$ (62)
@@
$T\left(\omega\right)\geq\sec h^{2}\left(\int_{r_{h}}^{+\infty}\frac{V_{s}}{2\omega}dr_{\ast}\right).$ (63)
@@
VI Energy Emission Rate: Energy Flux
@@
In the geometric-optics (high-frequency) regime, the absorption cross section oscillates around a constant limiting value $\sigma_{\rm lim}$. Motivated by the fact that the capture of high-energy quanta is governed by null geodesics, the shadow cast by the black hole causes the high energy cross section of absorption by the black hole. The limiting constant value of cross section area, which is related to the radius of the photon sphere is given as [55, 53, 93]
$\sigma_{\rm lim}\approx\pi R^{2}_{\rm sh},$ (65)
where $R_{\rm sh}$ denotes the radius of the black hole shadow.
Within this approximation, the spectral energy emission rate of black hole is given by the following equation [93]
$\frac{d^{2}\mathcal{E}}{d\omega\,dt}=\frac{2\pi^{2}\sigma_{\rm lim}}{e^{\omega/T}-1}\,\omega^{3},$ (66)
@@
The Hawking temperature of the black hole system is given by
$T=\frac{f^{\prime}(r_{h})}{4\pi}=\frac{1}{2\pi r_{h}}\left(\frac{M}{r_{h}}-\frac{Q^{2}}{r_{h}^{2}}\right),\quad r_{h}=M+\sqrt{M^{2}-Q^{2}}.$ (67)
As the selected space-time is asymptotically flat, the shadow radius of the black hole is given by
$R_{\rm sh}$ $=\frac{r_{s}}{\sqrt{f(r_{s})}}=\beta_{c}$
$=\frac{r_{s}}{\sqrt{\left(1-\frac{2M}{r_{s}}+\frac{Q^{2}}{r^{2}_{s}}\right)+\frac{\zeta^{2}}{r^{2}_{s}}\left(1-\frac{2M}{r_{s}}+\frac{Q^{2}}{r^{2}_{s}}\right)^{2}}},$ (68)
where $r_{s}$ is the photon sphere radius satisfying the following polynomial relation ($2\,f-r\,f^{\prime}=0$):
$r^{3}_{s}-3Mr^{2}_{s}+2Q^{2}r_{s}$
$+\zeta^{2}\left(r_{s}-5M+\frac{8M^{2}+4Q^{2}}{r_{s}}-\frac{12MQ^{2}}{r^{2}_{s}}+\frac{4Q^{4}}{r^{3}_{s}}\right)=0.$ (69)
@@
$\frac{d^{2}\mathcal{E}}{d\omega\,dt}=\frac{2\pi^{3}\omega^{3}\,r_{s}^{2}\,\left[\exp\left\{2\pi r_{h}\,\omega\,\left(\frac{M}{r_{h}}-\frac{Q^{2}}{r_{h}^{2}}\right)^{-1}\right\}-1\right]^{-1}}{\left(1-\frac{2M}{r_{s}}+\frac{Q^{2}}{r^{2}_{s}}\right)+\frac{\zeta^{2}}{r^{2}_{s}}\left(1-\frac{2M}{r_{s}}+\frac{Q^{2}}{r^{2}_{s}}\right)^{2}}.$ (70)
From the above analysis, we observe that the high-energy energy emission rate of the black hole is influenced by the quantum-correction parameter $\zeta$, the black hole mass $M$, and its electric charge $Q$.
(i) $\omega=0.1$
(ii) $\omega=0.3$
Figure 13: Behaviors of the energy emission rate given in Eq. (70) as a function of photon sphere radius $r_{s}$ by varying quantum-correction parameter $\zeta$. Here $Q/M=0.5$
Below, we discuss some special cases.
  • When $\zeta=0$, corresponding to the absence of quantum-correction in the black hole, the selected space-time simplifies to the Reissner-Nordstrom metric. In this limit, the energy flux simplifies as
    $\frac{d^{2}\mathcal{E}}{d\omega\,dt}=\frac{2\pi^{3}\omega^{3}\,r_{s}^{2}\,\left[\exp\left\{2\pi r_{h}\,\omega\,\left(\frac{M}{r_{h}}-\frac{Q^{2}}{r_{h}^{2}}\right)^{-1}\right\}-1\right]^{-1}}{\left(1-\frac{2M}{r_{s}}+\frac{Q^{2}}{r^{2}_{s}}\right)},$ (71)
    where $r_{h}$ is given earlier and $r_{s}$ is now as follows:
    $r_{s}=\frac{3M+\sqrt{9M^{2}-8Q^{2}}}{2},$ (72)
    the photon sphere radius for RN-black hole.
  • When $Q=0$, corresponding to the absence of the electric charge of the black hole, the selected space-time simplifies to the quantum-corrected black hole [98]. In that limiting case, the energy flux reduces to (setting now $r_{h}=2M$):
    $\frac{d^{2}\mathcal{E}}{d\omega\,dt}$ $=\frac{2\pi^{3}\omega^{3}\,r_{s}^{2}\,\left[e^{8\pi M\omega}-1\right]^{-1}}{\left(1-\frac{2M}{r_{s}}\right)+\frac{\zeta^{2}}{r^{2}_{s}}\left(1-\frac{2M}{r_{s}}\right)^{2}},$ (73)
    where $r_{s}$ here satisfies the following quartic equation:
    $r^{4}_{s}-3Mr^{3}_{s}+\zeta^{2}r_{s}^{2}-5M\zeta^{2}r_{s}+8M^{2}\zeta^{2}=0.$ (74)
    We observe that the energy flux is modified by the quantum-correction parameter $\zeta$, which consequently reduces the emission compared to the standard Schwarzschild black hole case.
(i) $\omega=0.1$
(ii) $\omega=0.3$
Figure 14: Behaviors of the energy emission rate given in Eq. (73) as a function of photon sphere radius $r_{s}$ by varying quantum-correction parameter $\zeta$.